Assume that Schmidt Machinery Company had the standard costs reflected in Exhibit 14.5. In a given month, the company used 3,870 pounds of aluminum to manufacture 950 units. The company paid $29.40 per pound during the month to purchase aluminum. At the beginning of the month, the company had 70 pounds of aluminum on hand. At the end of the month, the company had only 50 pound:s of aluminum in its warehouse. Schmidt used 4,700 direct labor hours during the month, at an average cost of $42.40 per hour Required: Compute for the month the following variances: 1. The purchase-price variance for aluminum. Indicate whether this variance is favorable (F) or unfavorable (U) 2. The usage variance for aluminum. Indicate whether this variance is favorable (F) or unfavorable (U). 3. The direct labor rate variance. Indicate whether this variance is favorable (F) ог unfavorable ( 4. The direct labor efficiency variance. Indicate whether this variance is favorable (F) or unfavorable (U) 1. Purchase-price variance 2. Usage variance 3. Direct labor rate variance 4. Efficiency variance
Standard Cost Sheet, Product XV-1 EXHIBIT 14.5 SCHMIDT MACHINERY COMPANY Standard Cost Sheet Product: XV-1 Descriptions Quantity Cost Rate Subtotal Total Direct materials: 4 pounds $25/pound $100 Aluminum 40 PVC I pound 40/pound 40/hour Direct labor 5 hours 200 Variable factory overhead 5 hours 12/hour Total variable manufacturing cost $400 24/hour 120 12 5 hours Fixed factory overhead Standard manufacturing cost per unit $520 50 Standard variable selling and administrative cost per unit "Budgeted fixed factory overhead cost $120,000
Answer #1

Solution 1:

Standard quantity of material (Aluminium) for actual production = 950*4 = 3800 Pound

Actual quantity of material (Aluminium) purchased = Quantity of material used + Ending inventory - Beginning inventory

= 3870 + 50 -70 = 3850 pound

Actual quantity of material (Aluminium) used = 3870 pound

Standard price of Aluminium = $25 per pound

Actual price of Aluminium = $29.40 per pound

Purchase price variance (Aluminium) = (SP - AP) * AQ = ($25 - $29.40) * 3850 = $16,940 U

Usage variance (Aluminium) = (SQ - AQ) * SR = (3800 - 3870) * $25 = $1,750 U

Solution 2:

Standard hours of direct labor = 950 * 5 = 4750 hours

Standard rate of direct labor = $40 per hour

Actual hours of direct labor = 4700 hours

Actual rate of direct labor =$42.40 per hour

Direct labor rate variance = (SR - AR) * AH = ($40 - $42.40) * 4700 = $11,280 U

Direct labor efficiency variance = (SH - AH) * SR = (4750- 4700) * $40 = $2,000 F